OpenStudy (anonymous):

Solve: 2x + 3y - 2z = -3 x + 9y - 2z = 3 3x + 6y + 2z = -1 I've tried pretty much every tactic I know to work this, can't figure it out. If someone could go step by step, that would help me out a lot.

OpenStudy (anonymous):

divide second eqn by 3 and divide third eqn by 2

OpenStudy (anonymous):

so u will get 3y in all eqns thn add em

OpenStudy (anonymous):

Substitution is the coolest! =D Pick your favorite equation and your favorite variable. Solve for that variable. Plug into either of the OTHER two equations. That gets you an equation with only 2 variables. Solve that equation for one of the remaining variables, plug into the 3rd equation.

OpenStudy (anonymous):

u can use Matrix to solve this, try it

OpenStudy (anonymous):

Smooth: That's what I did. But I came up with 5x + 9y = -4 and 4x + 15y = 2 and those two can't be multiplied in any form to just get x or y.

OpenStudy (anonymous):

hey linux have u tried the matrix approach , very simple & solvable

OpenStudy (anonymous):

Can't they though, Linux?

OpenStudy (anonymous):

I don't know what the matrix approach is. Lol.

OpenStudy (anonymous):

2x + 3y - 2z = -3 x + 9y - 2z = 3 3x + 6y + 2z = -1 solving the second equation for x: x = -9y +2z +3 substituting that into the first equation 3(-9y+2z+3)+3y-2z = -3 -24y + 4z =-12 Solving that for z gives z = -3 + 6y and plugging into the third equation for both x and z gives: 3x + 6y + 2z = -1 3(-9y +2z +3) + 6y +2(-3+6y) = -1 3(-9y +2(-3+6y)+3)+ 6y +2(-3+6y) = -1

OpenStudy (anonymous):

Before I continue, do you see what I've done so far?

OpenStudy (anonymous):

Yes, makes perfect sense now.

OpenStudy (anonymous):

That last step is especially tricky. You're sure you see how I substituted in for x and then for z twice?

OpenStudy (anonymous):

And that gives us a big messy equation, but it only has y in it, so we can solve it no problem.

OpenStudy (anonymous):

Yes, when I working it I was trying to solve for x or y, but it wasn't working right,

OpenStudy (anonymous):

Hmm. I'd love to see your work so that I could help you see your mistake.

OpenStudy (anonymous):

You should be able to solve for any variable you choose.

OpenStudy (anonymous):

Well, I took a and c first which gave me 5x + 9y = -4 and then I did b and c which gave me 4x + 15y = 2 then I tried went to put them together so I could change one and add them and then eventually solve for x, but I can't figure out how, since they can't be multipled in a way I can see.

OpenStudy (anonymous):

Ah, wait. I think I got the answers. x = -2 y = 2/3 and z = 1/2

OpenStudy (anonymous):

You can actually use those two equations, but it would require some fraction work. 5x+9y = -4 4x + 15y = 2 Solving equation 1 for x gives x = (-4/5) -(9/5)y substituting into equation 2 gives 4((-4/5)-(9/5)y) + 15y = 2 (-16/5)-(36/5)y + 15y = 2 (-16/5)+(39/5)y = 2 (39/5)y = 26/5 y = 26/39

OpenStudy (anonymous):

and 26/39 = 2/3

OpenStudy (anonymous):

So then my answers are correct?

OpenStudy (anonymous):

Yes! Good work =)